OpenStudy (anonymous):

5 times x6 square root

OpenStudy (anonymous):

\[5\sqrt{x ^{16}}\]

OpenStudy (anonymous):

5x^8?

OpenStudy (anonymous):

= x^(16/5)

OpenStudy (anonymous):

x^3.2

OpenStudy (anonymous):

yea i guess so if you want it written like that

OpenStudy (anonymous):

so that's the final answer

OpenStudy (anonymous):

what does the question say..

OpenStudy (anonymous):

simplify: \[5\sqrt{x ^{16}}\] answers: a)x^4 b) x^8 c) \[x ^{5\sqrt{x ^{3}}}\] d) x^3times 5 square root x

OpenStudy (anonymous):

oh so that is not the 5th root of 16? its 5 multiply by sqrt of (x^16)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):



OpenStudy (anonymous):

so its not like that?

OpenStudy (anonymous):

no it is 5 times square root of x to the 16th power

OpenStudy (anonymous):